OpenStudy (anonymous):

Solve. –6 ≥ 10 – 8x A. x ≥ –2 B. x ≤ –2 C. x ≥ 2 D. x ≤ 2

ganeshie8 (ganeshie8):

\(\large –6 \ge 10 – 8x\)

ganeshie8 (ganeshie8):

add 8x both sides

ganeshie8 (ganeshie8):

\(\large –6 \ge 10 – 8x\) \(\large –6+8x \ge 10 \)

ganeshie8 (ganeshie8):

next add 6 both sides

OpenStudy (anonymous):

2 and 18

ganeshie8 (ganeshie8):

hey its "8x" not just "8"

ganeshie8 (ganeshie8):

we cannot add apples and coconuts

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

so, -6 and 8x will stay like that, we dont add them away ok :)

OpenStudy (anonymous):

so its 2x and 18x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so C?

ganeshie8 (ganeshie8):

\(\large –6 \ge 10 – 8x\) add 8x both sides : \(\large –6+8x \ge 10 \) add 6 both sides \(\large 8x \ge 16\) divide 8 both sides \(\large x \ge 2\)

ganeshie8 (ganeshie8):

Yes !!

OpenStudy (anonymous):

tnx

ganeshie8 (ganeshie8):

u wlc :)